4.Your profit P is determined by subtracting the cost C (the amount of money it costs to operate a business) from the revenue R (the amount of money you earn from selling your product). Profit can be represented algebraically by the equations:
Profit=Revenue-Cost
OR
P = R - C
a) Rewrite the formula to solve for C.
b) Suppose your profit for one day is $1,281, and the cost of running the business for the day is $1,463. What is the revenue for that day? Explain your answer.

Answer by stanbon(75887)

Profit=Revenue-Cost
OR
P = R - C
a) Rewrite the formula to solve for C.
C = R-P
---------------
b) Suppose your profit for one day is $1,281, and the cost of running the business for the day is $1,463. What is the revenue for that day? Explain your answer.
R = C+P
R = 1463 + 1281
-------------------
R = 2744
================
Cheers,
Stan H.